OpenStudy (anonymous):

A polynomial has been factored, as shown below: f(x) = (x − 9)(x + 12)(x − 24) What are the zeros of the polynomial? −9, 9, −12, 12, −24, 24 −9, 12, −24 9, −12, 24 9, 12, 24

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

help

OpenStudy (anonymous):

i see you

OpenStudy (cggurumanjunath):

c

OpenStudy (anonymous):

how do i know its C?

OpenStudy (cggurumanjunath):

(x-9) =0

OpenStudy (cggurumanjunath):

(x+12)=0

OpenStudy (cggurumanjunath):

(x-24)=0

OpenStudy (cggurumanjunath):

equate each of those to zero ! and solve for it (x-24)=0 =>24

OpenStudy (cggurumanjunath):

now do u agree @lealealea222 ?

OpenStudy (anonymous):

yes thankyou

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

Anytime .